Mary is twice Wang's age & Kat is 2 years younger the Wang. The sum total of their ages is 46. How old is each person?
Found 2 solutions by anantha, MissionPossible:
Answer by anantha(86)

sol:
let Wang's age be x
Mary is twice Wnag's=2*x
kat is 2 years younger the wang=x-2
according to the problem
sum of their ages=46
x+2x+x-2=46
4x-2=46
adding 2 on both sides
4x-2+2=46+2
4x=48
dividing by 4 on both sides
4x/4=48/4
x=12
Wang's age=12 years
Mary age=2*Wang's age=2*12=24 years
Kat age=Wang's age-2=12-2=10 years
check:
sum of their ages=46
Wang's age+Mary age+Kat age=12+24+10=46
L.H.S=R.H.S

Answer by MissionPossible(18)

Mary is twice Wang's age & Kat is 2 years younger the Wang. The sum total of their ages is 46. How old is each person?
Mary's age = M
Wang's age = W
Kat's age = K
Mary is twice Wang's age - - > M = 2W
Kat is 2 years younger the Wang. - - -> K = W-2
Sum = 46 - - > M + W + K = 46
2W + W + (W-2) = 46
4W - 2 = 46
4W = 48
W = 48/4 = 12
M = 2W , then M = 24
K = W - 2 , then K = 12-2 = 10
Mary's age = 24 years
Wang's age = 12 years
Kat's age = 10 years